# Information That Matters: Exploring Information Needs of People Affected by Algorithmic Decisions

### v1

@@ 6. Limitations
Our participants were recruited from a local job agency and an apartment complex from the same geographical region, perhaps resulting in regional or cultural biases. Due to the limited sample size, we did not analyze the impact of sex and/or gender on our results. While this limits the generalizability of our results regarding sex and/or gender aspects, it does not limit the overall validity of our results. Further, the articulated information needs and observed processes of understanding are tied to the presented use cases, as demonstrated in Section 4.2, and to the verbal exchange with the study examiner, meaning that participants would likely ask different questions if information was presented in another fashion or by another person. Concerning the size of the participant sample, we are guided by research on qualitative methods, which suggests that 16 to 24 interviews are needed to reach adequate code and meaning saturation (Hennink et al., 2017).
@@ 7. Conclusion
The use of public AI systems in society can have considerable impact on people affected by algorithmic decisions. These affected stakeholders should be able to understand AI systems in order to take action, such as contesting decisions or asking for human intervention. Explanations can support this empowerment, but affected stakeholders’ information needs have rarely been covered in explanation approaches. To address this gap, we provide a collection of affected stakeholders’ information needs titled the ”XAI Novice Question Bank” (Figure 2), compiled from interviews with 24 participants and covering two ADM use cases (employment prediction and a health wristband). Our approach aims to extend the concept of question-driven explanations, inspired by the XAI Question Bank (Liao et al., 2020), to lay audiences and highlights the relevance of information needs about an ADM system’s context and usage. Further, we analyze participant understanding by applying cognitive theories to examine from which perspective participants approach information acquisition. We provide suggestions on matching explanations to the way stakeholders make sense of information and on using question-driven explanations to overcome understanding challenges. Lastly, we examine participants’ perceptions of the systems’ risk and benefits, finding that higher perceived risk likely prompts participants to inquire about intentional information, such as the motivation and consequences of deploying a system. We close with six key implications that our findings have for the design of explanations for affected stakeholders and posit that more research is needed to create suitable explanations for this crucial stakeholder group.
@@ Acknowledgements.
- Ackerman (2000) Mark S. Ackerman. 2000. The Intellectual Challenge of CSCW: The Gap Between Social Requirements and Technical Feasibility. Human–Computer Interaction 15, 2-3 (2000), 179–203. https://doi.org/10.1207/S15327051HCI1523_5
@@
- Alfrink et al. (2022) Kars Alfrink, Ianus Keller, Gerd Kortuem, and Neelke Doorn. 2022. Contestable AI by Design: Towards a Framework. Minds and Machines (13 Aug 2022). https://doi.org/10.1007/s11023-022-09611-z
- Ali et al. (2023) Sajid Ali, Tamer Abuhmed, Shaker El-Sappagh, Khan Muhammad, Jose M. Alonso-Moral, Roberto Confalonieri, Riccardo Guidotti, Javier Del Ser, Natalia Díaz-Rodríguez, and Francisco Herrera. 2023. Explainable Artificial Intelligence (XAI): What we know and what is left to attain Trustworthy Artificial Intelligence. Information Fusion 99 (Nov. 2023), 101805. https://doi.org/10.1016/j.inffus.2023.101805
- Allhutter (2021) Doris Allhutter. 2021. Ein Algorithmus zur effizienten Förderung der Chancen auf dem Arbeitsmarkt? WISO – Zeitschrift für Sozial- und Wirtschaftswissenschaften 44.JG (2021), 82–95.
@@
- Allhutter et al. (2020b) Doris Allhutter, Astrid Mager, Florian Cech, Fabian Fischer, and Gabriel Grill. 2020b. DER AMS-ALGORITHMUS: Eine Soziotechnische Analyse des Arbeitsmarktchancen-Assistenz-Systems (AMAS). Technical Report. https://epub.oeaw.ac.at/0xc1aa5576_0x003bfdf3.pdf
@@
- Baeza-Yate and Matthews (2022) Ricardo Baeza-Yate and Jeanna Matthews. 2022. Statement on Principles for Responsible Algorithmic Systems. https://www.acm.org/binaries/content/assets/public-policy/final-joint-ai-statement-update.pdf
@@
- Biran and Cotton (2017) Or Biran and Courtenay V. Cotton. 2017. Explanation and Justification in Machine Learning : A Survey Or.
@@
- Byrne (2023) Ruth M.J. Byrne. 2023. Good Explanations in Explainable Artificial Intelligence (XAI): Evidence from Human Explanatory Reasoning. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence. International Joint Conferences on Artificial Intelligence Organization, Macau, SAR China, 6536–6544. https://doi.org/10.24963/ijcai.2023/733
@@
- Commission (2021) European Commission. 2021. Laying Down Harmonised Rules on Artificial Intelligence. https://digital-strategy.ec.europa.eu/en/library/proposal-regulation-laying-down-harmonised-rules-artificial-intelligence
@@
- Council (2023) Swiss Federal Council. 2023. Popular vote. https://www.admin.ch/gov/en/start/documentation/votes.html. Official website of the Swiss Federal Council describing the popular vote. Last accessed: 14.10.2023.
@@
- Duckworth (2001) Eleanor Duckworth (Ed.). 2001. ”Tell me more”: listening to learners explain. Teachers College Press, New York.
@@
- European Parliament. Directorate General for Parliamentary Research Services. (2019) European Parliament. Directorate General for Parliamentary Research Services. 2019. Understanding algorithmic decision-making: opportunities and challenges. Publications Office, LU. https://data.europa.eu/doi/10.2861/536131
- Ferrario and Loi (2022) Andrea Ferrario and Michele Loi. 2022. How Explainability Contributes to Trust in AI. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 1457–1466. https://doi.org/10.1145/3531146.3533202
- Floridi (2019) Luciano Floridi. 2019. Establishing the rules for building trustworthy AI. Nature Machine Intelligence 1, 6 (June 2019), 261–262. https://doi.org/10.1038/s42256-019-0055-y
- Freiesleben and König (2023) Timo Freiesleben and Gunnar König. 2023. Dear XAI Community, We Need to Talk!. In Explainable Artificial Intelligence, Luca Longo (Ed.). Springer Nature Switzerland, Cham, 48–65.
@@
- Hagendorff (2022) Thilo Hagendorff. 2022. Blind spots in AI ethics. AI and Ethics 2, 4 (Nov. 2022), 851–867. https://doi.org/10.1007/s43681-021-00122-8
@@
- Holl et al. (2018) Jürgen Holl, Günter Kernbeiß, and Michael Wagner-Pinter. 2018. Das AMS-Arbeitsmarktchancen-Modell. Technical concept. Synthesis Forschung, Vienna.
@@
- Kaun and Taranu (2020) Anne Kaun and Gabriela Taranu. 2020. Automating Society Report 2020 / Sweden. https://automatingsociety.algorithmwatch.org/report2020/sweden/
@@
- Keil (2006) Frank C. Keil. 2006. Explanation and Understanding. Annual Review of Psychology 57, 1 (2006), 227–254. https://doi.org/10.1146/annurev.psych.57.102904.190100 arXiv:https://doi.org/10.1146/annurev.psych.57.102904.190100 PMID: 16318595.
- Keil (2021) Frank C. Keil. 2021. The Challenges and Benefits of Mechanistic Explanation in Folk Scientific Understanding. In Advances in experimental philosophy of science (paperback edition ed.), Daniel A. Wilkenfeld and Richard Samuels (Eds.). Bloomsbury Academic, London New York Oxford New Delhi Sydney, 41–56.
@@
- Kim et al. (2024) Minjung Kim, Saebyeol Kim, Jinwoo Kim, Tae-Jin Song, and Yuyoung Kim. 2024. Do stakeholder needs differ? - Designing stakeholder-tailored Explainable Artificial Intelligence (XAI) interfaces. International Journal of Human-Computer Studies 181 (2024), 103160. https://doi.org/10.1016/j.ijhcs.2023.103160
- Kim et al. (2023) Sunnie S. Y. Kim, Elizabeth Anne Watkins, Olga Russakovsky, Ruth Fong, and Andrés Monroy-Hernández. 2023. ”Help Me Help the AI”: Understanding How Explainability Can Support Human-AI Interaction. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 250, 17 pages. https://doi.org/10.1145/3544548.3581001
@@
- Kutsarova (2020) Viktoriya Kutsarova. 2020. Managing Alarming Situations with Mobile Crowdsensing Systems and Wearable Devices. KTH Royal Institute of Technology. School of Electrical Engineering and Computer Science. https://kth.diva-portal.org/smash/record.jsf?pid=diva2%3A1465193&dswid=-5365
- Kuziemski and Misuraca (2020) Maciej Kuziemski and Gianluca Misuraca. 2020. AI governance in the public sector: Three tales from the frontiers of automated decision-making in democratic settings. Telecommunications Policy 44, 6 (2020), 101976. https://doi.org/10.1016/j.telpol.2020.101976 Artificial intelligence, economy and society.
@@
- Liao et al. (2021) Q. Vera Liao, Milena Pribić, Jaesik Han, Sarah Miller, and Daby Sow. 2021. Question-Driven Design Process for Explainable AI User Experiences. arXiv:2104.03483 [cs.HC]
@@
- Lopez (2019) Paola Lopez. 2019. Reinforcing Intersectional Inequality via the AMS Algorithm in Austria. (2019), 21.
@@
- Lupia (2016) Arthur Lupia. 2016. Uninformed: why people know so little about politics and what we can do about it. Oxford University Press, New York.
- Madaio et al. (2020) Michael A. Madaio, Luke Stark, Jennifer Wortman Vaughan, and Hanna Wallach. 2020. Co-Designing Checklists to Understand Organizational Challenges and Opportunities around Fairness in AI. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–14. https://doi.org/10.1145/3313831.3376445
@@
- Mueller et al. (2019) Shane T. Mueller, Robert R. Hoffman, William Clancey, Abigail Emrey, and Gary Klein. 2019. Explanation in Human-AI Systems: A Literature Meta-Review, Synopsis of Key Ideas and Publications, and Bibliography for Explainable AI. arXiv:1902.01876 [cs.AI]
- Nannini et al. (2023) Luca Nannini, Agathe Balayn, and Adam Leon Smith. 2023. Explainability in AI Policies: A Critical Review of Communications, Reports, Regulations, and Standards in the EU, US, and UK. In 2023 ACM Conference on Fairness, Accountability, and Transparency. ACM, Chicago IL USA, 1198–1212. https://doi.org/10.1145/3593013.3594074
@@
- on Artificial Intelligence (2019) High-Level Expert Group on Artificial Intelligence. 2019. Ethics guidelines for trustworthy AI. (April 2019). https://digital-strategy.ec.europa.eu/en/library/ethics-guidelines-trustworthy-ai
@@
- Panigutti et al. (2023) Cecilia Panigutti, Ronan Hamon, Isabelle Hupont, David Fernandez Llorca, Delia Fano Yela, Henrik Junklewitz, Salvatore Scalzo, Gabriele Mazzini, Ignacio Sanchez, Josep Soler Garrido, and Emilia Gomez. 2023. The Role of Explainable AI in the Context of the AI Act. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 1139–1150. https://doi.org/10.1145/3593013.3594069
- Rozenblit and Keil (2002) Leonid Rozenblit and Frank Keil. 2002. The misunderstood limits of folk science: an illusion of explanatory depth. Cognitive Science 26, 5 (2002), 521–562. https://doi.org/10.1207/s15516709cog2605_1 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1207/s15516709cog2605_1
@@
- Sahlen et al. (2018) Malin Sahlen, Edvin Alam, and Paulo Silva. 2018. Ny teknik i välfärden - vår tids välfärdskliv (New Technologies in Welfare – The Welfare Leap of Our Times). Technical Report. Svenskt Näringsliv. https://www.svensktnaringsliv.se/sakomraden/valfard-och-offentlig-sektor/ny-teknik-i-valfarden-var-tids-valfardskliv_1003696.html
@@
- Schmude et al. (2023a) Timothée Schmude, Laura Koesten, Torsten Möller, and Sebastian Tschiatschek. 2023a. On the Impact of Explanations on Understanding of Algorithmic Decision-Making. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 959–970. https://doi.org/10.1145/3593013.3594054
- Schmude et al. (2023b) Timothée Schmude, Laura Koesten, Torsten Möller, and Sebastian Tschiatschek. 2023b. Applying Interdisciplinary Frameworks to Understand Algorithmic Decision-Making. https://doi.org/10.48550/arXiv.2305.16700 arXiv:2305.16700 [cs.HC]
- Schoeffer et al. (2022) Jakob Schoeffer, Niklas Kuehl, and Yvette Machowski. 2022. “There Is Not Enough Information”: On the Effects of Explanations on Perceptions of Informational Fairness and Trustworthiness in Automated Decision-Making. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 1616–1628. https://doi.org/10.1145/3531146.3533218
- Scott et al. (2022) Kristen M. Scott, Sonja Mei Wang, Milagros Miceli, Pieter Delobelle, Karolina Sztandar-Sztanderska, and Bettina Berendt. 2022. Algorithmic Tools in Public Employment Services: Towards a Jobseeker-Centric Perspective. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 2138–2148. https://doi.org/10.1145/3531146.3534631
- Shen et al. (2020) Hong Shen, Haojian Jin, Ángel Alexander Cabrera, Adam Perer, Haiyi Zhu, and Jason I. Hong. 2020. Designing Alternative Representations of Confusion Matrices to Support Non-Expert Public Understanding of Algorithm Performance. Proceedings of the ACM on Human-Computer Interaction 4, CSCW2 (Oct. 2020), 1–22. https://doi.org/10.1145/3415224
@@
- Sovrano et al. (2022) Francesco Sovrano, Salvatore Sapienza, Monica Palmirani, and Fabio Vitali. 2022. Metrics, Explainability and the European AI Act Proposal. J 5, 1 (2022), 126–138. https://doi.org/10.3390/j5010010
@@
- Thiebes et al. (2021) Scott Thiebes, Sebastian Lins, and Ali Sunyaev. 2021. Trustworthy Artificial Intelligence. Electronic Markets 31, 2 (June 2021), 447–464. https://doi.org/10.1007/s12525-020-00441-4
- Veale et al. (2018) Michael Veale, Max Van Kleek, and Reuben Binns. 2018. Fairness and Accountability Design Needs for Algorithmic Support in High-Stakes Public Sector Decision-Making. Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (April 2018), 1–14. https://doi.org/10.1145/3173574.3174014
- Veale and Zuiderveen Borgesius (2021) Michael Veale and Frederik Zuiderveen Borgesius. 2021. Demystifying the Draft EU Artificial Intelligence Act. preprint. SocArXiv. https://doi.org/10.31235/osf.io/38p5f
- Vredenburgh (2022) Kate Vredenburgh. 2022. The Right to Explanation*. Journal of Political Philosophy 30, 2 (June 2022), 209–229. https://doi.org/10.1111/jopp.12262
@@
- Wiggins and McTighe (2005) Grant P. Wiggins and Jay McTighe. 2005. Understanding by design (expanded 2nd ed ed.). Association for Supervision and Curriculum Development, Alexandria, VA. OCLC: 60756429.

### v6

@@ 6. Limitations
Our participants were recruited from a local job agency and an apartment complex from the same geographical region, perhaps resulting in regional or cultural biases. Due to the limited sample size, we did not analyze the impact of sex and/or gender on our results. While this limits the generalizability of our results regarding sex and/or gender aspects, it does not limit the overall validity of our results. Further, the articulated information needs and observed processes of understanding are tied to the presented use cases, as demonstrated in Section 4.2, and to the verbal exchange with the study examiner, meaning that participants would likely ask different questions if the information was presented in another fashion or by another person. Concerning the size of the participant sample, we are guided by research on qualitative methods, which suggests that 16 to 24 interviews are needed to reach adequate code and meaning saturation (Hennink et al., 2017).
@@ 7. Conclusion
The use of public AI systems in society can have a considerable impact on people affected by algorithmic decisions. These stakeholders should be able to understand AI systems in order to take action, such as contesting decisions or asking for human intervention. Explanations can support this empowerment, but the information needs of this stakeholder group—who tend to be AI novices—have rarely been covered in explanation approaches. To address this gap, we provide a collection of information needs titled the “XAI Novice Question Bank” (Figure 3), compiled from interviews with 24 participants and covering two ADM use cases (employment prediction and a health wristband). Our approach aims to extend the concept of question-driven explanations, inspired by the XAI Question Bank (Liao et al., 2020), to lay audiences and highlights the relevance of information needs about an ADM system’s context and usage. Further, we analyze participant understanding by applying cognitive theories to examine from which perspective participants approach information acquisition. We provide suggestions on matching explanations to how stakeholders make sense of information and strategies to overcome understanding challenges. Lastly, we examine participants’ perceptions of the systems’ risk and benefits, finding that higher perceived risk likely prompts participants to inquire about intentional information, such as the motivation and consequences of deploying a system. We close with five critical implications that our findings have for the design of explanations for AI novices affected by algorithmic decisions. More research is needed to create suitable explanations for this crucial stakeholder group.
@@ Acknowledgements.
- Alfrink et al. (2023) Kars Alfrink, Ianus Keller, Gerd Kortuem, and Neelke Doorn. 2023. Contestable AI by Design: Towards a Framework. Minds and Machines 33, 4 (01 Dec 2023), 613–639. https://doi.org/10.1007/s11023-022-09611-z
@@
- Allhutter et al. (2020b) Doris Allhutter, Astrid Mager, Florian Cech, Fabian Fischer, and Gabriel Grill. 2020b. Der AMS-Algorithmus: Eine Soziotechnische Analyse des Arbeitsmarktchancen-Assistenz-Systems (AMAS). Technical Report. epub.oeaw.ac.at.
@@
- Bertrand et al. (2023) Astrid Bertrand, James R. Eagan, and Winston Maxwell. 2023. Questioning the ability of feature-based explanations to empower non-experts in robo-advised financial decision-making. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (¡conf-loc¿, ¡city¿Chicago¡/city¿, ¡state¿IL¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 943–958. https://doi.org/10.1145/3593013.3594053
- Biran and Cotton (2017) Or Biran and Courtenay V. Cotton. 2017. Explanation and Justification in Machine Learning : A Survey. In Proceedings of the IJCAI 2017 Workshop on Explainable Artificial Intelligence (XAI). cs.columbia.edu/ orb/papers/xai_survey_paper_2017.pdf.
@@
- Bove et al. (2022) Clara Bove, Jonathan Aigrain, Marie-Jeanne Lesot, Charles Tijus, and Marcin Detyniecki. 2022. Contextualization and Exploration of Local Feature Importance Explanations to Improve Understanding and Satisfaction of Non-Expert Users. In Proceedings of the 27th International Conference on Intelligent User Interfaces (¡conf-loc¿, ¡city¿Helsinki¡/city¿, ¡country¿Finland¡/country¿, ¡/conf-loc¿) (IUI ’22). Association for Computing Machinery, New York, NY, USA, 807–819. https://doi.org/10.1145/3490099.3511139
@@
- Bundesagentur für Arbeit (2021) Bundesagentur für Arbeit. 2021. Bearbeiten von Bewerberdaten durch Träger. arbeitsagentur.de/datei/dok_ba013193.pdf.
@@
- Byrne (2023) Ruth M.J. Byrne. 2023. Good Explanations in Explainable Artificial Intelligence (XAI): Evidence from Human Explanatory Reasoning. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence. Macau, SAR China, 6536–6544. https://doi.org/10.24963/ijcai.2023/733
@@
- Cavaliere and Romeo (2022) Paolo Cavaliere and Graziella Romeo. 2022. From Poisons to Antidotes: Algorithms as Democracy Boosters. European Journal of Risk Regulation 13, 3 (Sept. 2022), 421–442. https://doi.org/10.1017/err.2021.57
@@
- Corvite et al. (2023) Shanley Corvite, Kat Roemmich, Tillie Ilana Rosenberg, and Nazanin Andalibi. 2023. Data Subjects’ Perspectives on Emotion Artificial Intelligence Use in the Workplace: A Relational Ethics Lens. Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (April 2023), 1–38. https://doi.org/10.1145/3579600
@@
- Desiere and Struyven (2021) Sam Desiere and Ludo Struyven. 2021. Using Artificial Intelligence to classify Jobseekers: The Accuracy-Equity Trade-off. Journal of Social Policy 50, 2 (2021), 367–385. https://doi.org/10.1017/S0047279420000203
@@
- Duckworth (2001) Eleanor Duckworth (Ed.). 2001. “Tell me more”: Listening to learners explain. Teachers College Press, New York.
@@
- European Commission (2021) European Commission. 2021. Laying Down Harmonised Rules on Artificial Intelligence. digital-strategy.ec.europa.eu/en/library/proposal-regulation-laying-down-harmonised-rules-artificial-intelligence.
- European Parliament. Directorate General for Parliamentary Research Services. (2019) European Parliament. Directorate General for Parliamentary Research Services. 2019. Understanding algorithmic decision-making: opportunities and challenges. Publications Office, LU. https://doi.org/10.2861/536131
- Freiesleben and König (2023) Timo Freiesleben and Gunnar König. 2023. Dear XAI Community, We Need to Talk!. In Explainable Artificial Intelligence, Luca Longo (Ed.). Springer Nature Switzerland, Cham, 48–65. https://doi.org/10.1007/978-3-031-44064-9_3
- Gerke et al. (2020) Sara Gerke, Timo Minssen, and Glenn Cohen. 2020. Chapter 12 - Ethical and legal challenges of artificial intelligence-driven healthcare. In Artificial Intelligence in Healthcare, Adam Bohr and Kaveh Memarzadeh (Eds.). Academic Press, 295–336. https://doi.org/10.1016/B978-0-12-818438-7.00012-5
@@
- Holl et al. (2018) Jürgen Holl, Günter Kernbeiß, and Michael Wagner-Pinter. 2018. Das AMS-Arbeitsmarktchancen-Modell.
- Holl et al. (2019) Jürgen Holl, Günter Kernbeiß, and Michael Wagner-Pinter. 2019. Personenbezogene Wahrscheinlichkeitsaussagen (»Algorithmen«): Stichworte zur Sozialverträglichkeit. Technical concept. Synthesis Forschung, Vienna.
@@
- Karinshak et al. (2023) Elise Karinshak, Sunny Xun Liu, Joon Sung Park, and Jeffrey T. Hancock. 2023. Working With AI to Persuade: Examining a Large Language Model’s Ability to Generate Pro-Vaccination Messages. Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (April 2023), 1–29. https://doi.org/10.1145/3579592
- Kaun and Taranu (2020) Anne Kaun and Gabriela Taranu. 2020. Automating Society Report 2020 / Sweden. automatingsociety.algorithmwatch.org/report2020/sweden/.
@@
- Keil (2006) Frank C. Keil. 2006. Explanation and Understanding. Annual Review of Psychology 57, 1 (2006), 227–254. https://doi.org/10.1146/annurev.psych.57.102904.190100
- Keil (2021) Frank C. Keil. 2021. The Challenges and Benefits of Mechanistic Explanation in Folk Scientific Understanding. In Advances in experimental philosophy of science (paperback ed.), Daniel A. Wilkenfeld and Richard Samuels (Eds.). Bloomsbury Academic, London New York Oxford New Delhi Sydney, 41–56.
@@
- Kim et al. (2024) Minjung Kim, Saebyeol Kim, Jinwoo Kim, Tae-Jin Song, and Yuyoung Kim. 2024. Do stakeholder needs differ? Designing stakeholder-tailored Explainable Artificial Intelligence (XAI) interfaces. International Journal of Human-Computer Studies 181 (2024), 103160. https://doi.org/10.1016/j.ijhcs.2023.103160
- Kim et al. (2023) Sunnie S. Y. Kim, Elizabeth Anne Watkins, Olga Russakovsky, Ruth Fong, and Andrés Monroy-Hernández. 2023. “Help Me Help the AI”: Understanding How Explainability Can Support Human-AI Interaction. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 250, 17 pages. https://doi.org/10.1145/3544548.3581001
@@
- Kutsarova (2020) Viktoriya Kutsarova. 2020. Managing Alarming Situations with Mobile Crowdsensing Systems and Wearable Devices. Master’s thesis. KTH, School of Electrical Engineering and Computer Science (EECS).
- Kuziemski and Misuraca (2020) Maciej Kuziemski and Gianluca Misuraca. 2020. AI governance in the public sector: Three tales from the frontiers of automated decision-making in democratic settings. Telecommunications Policy 44, 6 (2020), 101976. https://doi.org/10.1016/j.telpol.2020.101976
@@
- Liao et al. (2021) Q. Vera Liao, Milena Pribić, Jaesik Han, Sarah Miller, and Daby Sow. 2021. Question-Driven Design Process for Explainable AI User Experiences. arXiv:https://arxiv.org/abs/2104.03483 [cs.HC]
@@
- Lopez (2019) Paola Lopez. 2019. Reinforcing Intersectional Inequality via the AMS Algorithm in Austria. In Proceedings of the 18th Annual STS Conference. Graz, 289–309. https://doi.org/10.3217/978-3-85125-668-0-16
@@
- Lupia (2016) Arthur Lupia. 2016. Uninformed: Why people know so little about politics and what we can do about it. Oxford University Press, New York. https://doi.org/10.1093/oso/9780190263720.001.0001
- Martijn et al. (2022) Millecamp Martijn, Cristina Conati, and Katrien Verbert. 2022. “Knowing me, knowing you”: personalized explanations for a music recommender system. User Modeling and User-Adapted Interaction 32, 1-2 (April 2022), 215–252. https://doi.org/10.1007/s11257-021-09304-9
@@
- Mueller et al. (2019) Shane T. Mueller, Robert R. Hoffman, William Clancey, Abigail Emrey, and Gary Klein. 2019. Explanation in Human-AI Systems: A Literature Meta-Review, Synopsis of Key Ideas and Publications, and Bibliography for Explainable AI. arXiv:https://arxiv.org/abs/1902.01876
- Nahavandi et al. (2022) Darius Nahavandi, Roohallah Alizadehsani, Abbas Khosravi, and U Rajendra Acharya. 2022. Application of artificial intelligence in wearable devices: Opportunities and challenges. Computer methods and programs in biomedicine 213 (2022), 106541–106541. https://doi.org/10.1016/j.cmpb.2021.106541
@@
- Phillips et al. (2021) P Jonathon Phillips, Carina A Hahn, Peter C Fontana, Amy N Yates, Kristen Greene, David A Broniatowski, and Mark A Przybocki. 2021. Four principles of explainable artificial intelligence. Technical Report NIST IR 8312. National Institute of Standards and Technology (U.S.), Gaithersburg, MD. NIST IR 8312 pages. https://doi.org/10.6028/NIST.IR.8312
- Raji et al. (2022) Inioluwa Deborah Raji, I. Elizabeth Kumar, Aaron Horowitz, and Andrew Selbst. 2022. The Fallacy of AI Functionality. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 959–972. https://doi.org/10.1145/3531146.3533158
- Ram (1989) Ashwin Ram. 1989. Question-driven understanding: An integrated theory of story understanding, memory and learning. Ph. D. Dissertation. USA. Advisor(s) Schank, Roger C. https://doi.org/10.5555/916448
- Rozenblit and Keil (2002) Leonid Rozenblit and Frank Keil. 2002. The misunderstood limits of folk science: An illusion of explanatory depth. Cognitive Science 26, 5 (2002), 521–562. https://doi.org/10.1207/s15516709cog2605_1
- Rudin (2019) Cynthia Rudin. 2019. Stop explaining black box machine learning models for high stakes decisions and use interpretable models instead. Nature Machine Intelligence 1, 5 (01 May 2019), 206–215. https://doi.org/10.1038/s42256-019-0048-x
@@
- Sahlen et al. (2018) Malin Sahlen, Edvin Alam, and Paulo Silva. 2018. Ny teknik i välfärden - vår tids välfärdskliv (New Technologies in Welfare – The Welfare Leap of Our Times). Technical Report. Svenskt Näringsliv. svensktnaringsliv.se/sakomraden.
@@
- Schmude et al. (2023) Timothée Schmude, Laura Koesten, Torsten Möller, and Sebastian Tschiatschek. 2023. On the Impact of Explanations on Understanding of Algorithmic Decision-Making. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 959–970. https://doi.org/10.1145/3593013.3594054
- Schoeffer et al. (2022) Jakob Schoeffer, Niklas Kuehl, and Yvette Machowski. 2022. “There Is Not Enough Information”: On the Effects of Explanations on Perceptions of Informational Fairness and Trustworthiness in Automated Decision-Making. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 1616–1628. https://doi.org/10.1145/3531146.3533218
- Scott et al. (2022) Kristen M. Scott, Sonja Mei Wang, Milagros Miceli, Pieter Delobelle, Karolina Sztandar-Sztanderska, and Bettina Berendt. 2022. Algorithmic Tools in Public Employment Services: Towards a Jobseeker-Centric Perspective. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). ACM, Seoul Republic of Korea, 2138–2148. https://doi.org/10.1145/3531146.3534631
- Seneviratne et al. (2017) Suranga Seneviratne, Yining Hu, Tham Nguyen, Guohao Lan, Sara Khalifa, Kanchana Thilakarathna, Mahbub Hassan, and Aruna Seneviratne. 2017. A Survey of Wearable Devices and Challenges. IEEE Communications Surveys & Tutorials PP (07 2017), 1–1. https://doi.org/10.1109/COMST.2017.2731979
- Shen et al. (2020) Hong Shen, Haojian Jin, Angel Alexander Cabrera, Adam Perer, Haiyi Zhu, and Jason I. Hong. 2020. Designing Alternative Representations of Confusion Matrices to Support Non-Expert Public Understanding of Algorithm Performance. Proceedings of the ACM on Human-Computer Interaction 4, CSCW2 (Oct. 2020), 1–22. https://doi.org/10.1145/3415224
@@
- Shin (2022) Donghee Shin. 2022. How do people judge the credibility of algorithmic sources? AI & SOCIETY 37, 1 (March 2022), 81–96. https://doi.org/10.1007/s00146-021-01158-4
- Shin et al. (2022a) Donghee Shin, Kerk F. Kee, and Emily Y. Shin. 2022a. Algorithm awareness: Why user awareness is critical for personal privacy in the adoption of algorithmic platforms? International Journal of Information Management 65 (Aug. 2022), 102494. https://doi.org/10.1016/j.ijinfomgt.2022.102494
- Shin et al. (2022b) Donghee Shin, Joon Soo Lim, Norita Ahmad, and Mohammed Ibahrine. 2022b. Understanding user sensemaking in fairness and transparency in algorithms: Algorithmic sensemaking in over-the-top platform. AI & SOCIETY 39, 2 (April 2022), 477–490. https://doi.org/10.1007/s00146-022-01525-9
- Shin and Shin (2023) Donghee Shin and Emily Y. Shin. 2023. Human-Centered AI: A Framework for Green and Sustainable AI. Computer 56, 6 (2023), 16–25. https://doi.org/10.1109/MC.2023.3241071
- Shin (2023) Don Donghee Shin. 2023. Algorithms, humans, and interactions: How do algorithms interact with people? designing meaningful AI experiences (first edition ed.). Routledge, Boca Raton, FL. https://doi.org/10.1201/b23083
- Shneiderman (2022) Ben Shneiderman. 2022. Human-centered AI. Oxford University Press, Oxford.
@@
- Swiss Federal Council (2023) Swiss Federal Council. 2023. Popular vote. admin.ch/gov/en/start/documentation/votes.html.
@@
- Urquhart et al. (2009) Cathy Urquhart, Hans Lehmann, and Michael D. Myers. 2009. Putting the ‘theory’ back into grounded theory: guidelines for grounded theory studies in information systems: Guidelines for grounded theory studies in information systems. Information Systems Journal 20, 4 (May 2009), 357–381. https://doi.org/10.1111/j.1365-2575.2009.00328.x
@@
- Weiss (1995) Robert Stuart Weiss. 1995. Learning from strangers: The art and method of qualitative interview studies (1. free press paperback ed.). Free Press, New York, NY.
- Wieringa (2023) Maranke Wieringa. 2023. “Hey SyRI, tell me about algorithmic accountability”: Lessons from a landmark case. Data & Policy 5 (2023), e2. https://doi.org/10.1017/dap.2022.39
- Wiggins and McTighe (2005) Grant P. Wiggins and Jay McTighe. 2005. Understanding by design (expanded 2nd ed.). Association for Supervision and Curriculum Development, Alexandria, VA.